OpenStudy (anonymous):

Find all solutions x^2-4i=0 WILL METAL

OpenStudy (anonymous):

(x-?)(x+?) whats the ?=

OpenStudy (anonymous):

\[\pm2i\] =your answer

OpenStudy (anonymous):

\[x^2-4i=(x-2\sqrt i)(x+2\sqrt i)=0\]